OpenStudy (rsadhvika):

If a sequence \(\{a_n\}\) converges to both \(m\) and \(n\), prove that \(m = n\)

OpenStudy (anonymous):

Assume \(m\not=n\). If \(\{a_n\}\) converges to both of these values, then there exist \(N_1\) and \(N_2\) such that \(m\ge N_1\) and \(n\ge N_2\) imply, respectively, \[|a_n-m|<\epsilon_1=\frac{\epsilon}{2}\quad\text{and}\quad|a_n-n|<\epsilon_2=\frac{\epsilon}{2}\] for a given \(\epsilon>0\). (The fact that \(\epsilon_1=\epsilon_2\) is due to \(\epsilon\) being arbitrary.) What does this tell you?

OpenStudy (rsadhvika):

that tells \(N_1\) is with in \(\epsilon/2\) of\(m\) and \(N_2\) is with in \(\epsilon/2\) of \(m\) ?

OpenStudy (rsadhvika):

and sorry for mixing index variable with the limit, they are supposed to be different shall i change the index variable to i ?

OpenStudy (anonymous):

It doesn't matter to me :P Do you see where this proof is going? We assume there are two distinct limits, but we want to show that this leads to a contradiction.

OpenStudy (rsadhvika):

it would be contradiction for a sequence to converge to two limits, im struggling with putting steps

OpenStudy (rsadhvika):

** that tells \(N_1\) is with in \(\epsilon/2\) of \(m\) and \(N_2\) is with in \(\epsilon/2\) of \(n\)

OpenStudy (anonymous):

The contradiction here would be to show that if \(m\not=n\), we still arrive at \(m=n\). To do this, we could consider it in terms of the distance between these "distinct" limits. Since \(\epsilon>0\) can be arbitrarily small, if you can show that \(|m-n|<\epsilon\), then you are finished.

OpenStudy (rsadhvika):

Okay I see \[|a_n-m|< \frac{\epsilon}{2}\quad\text{and}\quad|a_n-n|<\frac{\epsilon}{2}\]

OpenStudy (rsadhvika):

what would be my next step ?

OpenStudy (anonymous):

How can you combine the absolute value terms?

OpenStudy (anonymous):

If we add corresponding sides together, we'll arrive at something close to what we want: \[|a_n-m|+|a_n-n|<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon\]

OpenStudy (rsadhvika):

the indexes are different right ?

OpenStudy (rsadhvika):

i mean the first absolute value inequality holds only when n > N1 and the second absolute value inrquality holds only when n > N2

OpenStudy (anonymous):

Sure, but if we only consider the greatest of the two, i.e. \(N=\max\{N_1,N_2\}\), then we don't have to worry about that.

OpenStudy (rsadhvika):

Ohk that makes sense : n > max(N1, N2)

OpenStudy (rsadhvika):